The rate of heat flow through a slab is H. If the slab thickness is halved, its cross-sectional area is doubled, and the temperature difference across it is doubled, then the rate of heat flow becomes (please show work)
Answer #1

Heat flow rate is given by:

H = k*A*dT/dx

k = thermal conductivity of slab

A = Cross-sectional Area of slab

dT = temperature difference across slab

dx = thickness of slab

Since 'k' for slab is constant, So

H2/H1 = (A2/A1)*(dT2/dT1)*(dx1/dx2)

Given that A2/A1 = 2

dT2/dT1 = 2

And dx2/dx1 = 1/2, So dx1/dx2 = 2

So,

H2/H1 = 2*2*2

H2 = 8*H1

So rate of heat flow across slab is increase by a factor of 8

Let me know if you've any query.
Answer #2

Solution: Rate of Heat Flow Under Modified Conditions

Given:

  • Initial rate of heat flow (H):

    H=kAΔTd

    where:

    • k = thermal conductivity (constant),

    • A = cross-sectional area,

    • ΔT = temperature difference,

    • d = thickness of the slab.


Modified Conditions:

  1. Thickness (d): Halved → d=d2.

  2. Area (A): Doubled → A=2A.

  3. Temperature difference (ΔT): Doubled → ΔT=2ΔT.


New Rate of Heat Flow (H):

H=kAΔTd=k(2A)(2ΔT)d/2=4kAΔTd/2=8(kAΔTd)=8H


Answer:

The new rate of heat flow becomes 8H.


answered by: anonymous